# SimuCourt: Building Judicial Decision-Making Agents with Real-world Judgement Documents

### v1

@@ 3.4 Judicial Knowledge Base Construction
Journal Articles Journal articles, typically authored by legal experts, can provide in-depth analysis and unique perspectives on specific legal issues. We collect highly-cited journal articles from 2010 to 2023 from the Chinese Legal Resources Knowledge Database  [^4] https://lawnew.cnki.net/. These articles span various legal fields, including but not limited to civil law, commercial law, criminal law and administrative law.
Precedents We collect all judgement documents of criminal, civil and administrative cases from the China Judgements Online for the years 2017 to 2022. However, as illustrated in Figure 7 in the Appendix, the data exhibits a significant long-tail distribution. To balance the type of case, we limit the number of cases for each cause of action to no more than 20k. For those causes of action with more cases, we retain only the top 20k cases with the longest text as representatives of complex cases. This comprehensive and diverse repository of precedents facilitates the identification of similar cases and provides effective references and guidance for judges when dealing with new cases.
@@
Table 2: Statistics of our judicial knowledge base.
3.5 Task Formulation
We propose a generative task to evaluate agent as judge. Specifically, as shown in Figure 1, we formulate the Judicial Decision-Making task as given the case details of a case, such as Determine facts, Complaint/Indictment, Statement of the plaintiff and the defendant, the agent system needs to make a complete judicial decision, which includes a clear and reasonable case analysis, rigorous legal grounds, and definitive final judgement. SimuCourt encompasses two experimental settings:
First Instance This setting refers to the trial court level, where the judge listens to arguments, determines the guilt of the defendant, and assesses whether punitive measures are warranted. Within this setting, the primary focus is on evaluating the agent’s understanding of relevant laws and its analysis of case facts.
Second Instance This setting refers to the appellate court level. During this stage, the judge re-evaluates the case, considering new evidence. The objective at this stage is to ensure the legality and fairness of the initial judgement, identifying legal errors or inappropriate application of regulations from the first instance and demonstrating the capability to effectively handle new evidence. Through these assessments, we aim to comprehensively evaluate the agent’s legal intelligence and logical reasoning abilities in judicial practice.
4 The AgentsCourt Framework
4.1 Court Debate Simulation
The court debate provides a platform for all parties involved to present their points and arguments comprehensively and fairly, which can significantly influence the judgement of the case. However, due to the majority of judgement documents only recording the key points of the plaintiff’s and defendant’s statements, obtaining complete court transcripts is challenging. Fortunately, as large language models have shown remarkable ability in role-playing Li et al. (2023); Qian et al. (2023); Chen et al. (2023), in this module, we aim to reconstruct the court debate with multiple agents for each case.
Simulated Process Since we have already collected determine facts of each case from judgement documents, we simplify the simulated court process into four stages: opening remarks; court debate; final statements; judicial decision. We set up three agents to play the roles of the judge, plaintiff, and defendant respectively. During the court session, the judge agent first delivers opening remarks, which include basic information about the plaintiff and the defendant, determination of facts, and so on. Then, the trial moves into the court debate stage and the communication between the agents will be recorded as court transcripts.
Court Debate In this stage, both the plaintiff and the defendant need to present their arguments in line with their interests. The plaintiff should vigorously argue their complaint, articulating their stance and reasoning. Meanwhile, the defendant must defend their actions, aiming to prove their innocence or seek a lighter penalty. For each agent, we carefully design an role-playing prompt to build their character personality and use the actual statements from judgment documents as the their starting prompts. It is worth noting that due to the limited record of statements in judgment documents, we combine the plaintiff and their representative, as well as the defendant and their representative, into the plaintiff and defendant, respectively, without setting separate roles for representatives.
4.2 Legal Information Retrieval
Court debate serves to thoroughly explore the facts and contentious issues within a case, making the judge better comprehend the complexity of the matter. Furthermore, to make accurate judicial decisions, judges must possess extensive legal information.
Figure 4: Automatic retrieval of precedents.
Automatic Information Retrieval In terms of knowledge base retrieval, as presented in Figure 4, the assistant first predict the type of case based on the determine facts of the current case. Due to the vast number of documents in the knowledge base, and the fact that cases with the same cause often have more similar keywords, we employs the BM25 model Lin et al. (2021) for efficient rough retrieval to obtain the top 100 documents from the knowledge base. Building on this, we further utilize the BGE-Large model Xiao et al. (2023) to encode and re-rank these retrieved documents and choose the most similar document to the current case as the optimal precedent. Additionally, to obtain more comprehensive laws and regulations relevant to the current case without introducing additional context, the judge assistant extracts the corresponding legal grounds from the top 5 precedents as related legal provisions of current case.
4.3 Judgement Refinement
Preliminary Judgement As shown in the bottom of Figure 3, after receiving the determine facts of current case and transcripts of simulated court debate, the judge agent takes the action of analysis, then provides its legal grounds and subsequently reaching a preliminary judgement. Here is the preliminary judgement of judge regarding the case in Figure 1:
Analysis: The court finds John Doe premeditatedly stole significant assets, constituting theft. His criminal record and rapid reoffense warrant a stricter sentence ... Legal grounds: Articles 65, Paragraph 3 of Article 67, and Article 264 of the Criminal Law of the People’s Republic of China. Judgement: The accused is found guilty of theft and is hereby sentenced to three years and six months of fixed-term imprisonment, and fined 5,000 RMB.
@@ 5 Experiments
Model Legal Grounds Judgement Results Case Analysis
@@
5.1 Automatic Evaluation Metrics
As illustrated in Appendix D, the legal grounds and judgement are concise and structured. Therefore, we propose corresponding automatic evaluation metrics for each and we will make the evaluation code publicly available.
Legal Grounds Evaluation The correct legal grounds is crucial for a fair judgment. Thus, we employ the strict matching method to assess the legal grounds generated by the agent system. Specifically, we calculate the number of entries that match and do not match between the legal grounds list of the agent system and the reference legal grounds list. These counts are then micro-averaged to determine the overall precision, recall and F1 scores.
Judgement Evaluation for Civil and Administrative Cases The judgment of each civil or administrative case may encompass multiple results, such as the confirmation of legal obligations, compensation orders, and the allocation of litigation costs. While each result typically revolves around a single key point, it may involve specific monetary amounts and interest rate information. Consequently, traditional text matching methods based on similarity struggle to accurately capture these key points. Thus, we employ GPT-4 as an evaluator. We separately count the number of matching and non-matching key points in the agent system’s judgment results compared to the reference judgment results. The micro-averaged counts are used to calculate the overall precision, recall and F1 scores.
@@ 5.2 Human Evaluation
We employ manual assessment for the generated case analysis. This is because case analysis entails intricate logical reasoning and ethical considerations that are challenging to evaluate through automatic metrics or GPT-4. For each setting, we present a panel of three undergraduate students a random sample of 100 entries from each setting and the following binary True/False criteria guidelines: 1) Correctness: Mark true if and only if the analysis is satisfying and considers all parties involved. 2) Logicality: Mark false if the analysis contains any illogical or untrue reasoning. 3) Concision: Mark true if the analysis covers all necessary information without any extra information.
@@ 5.4 Main Results
As shown in Table 3, our framework outperforms other models in all aspects. For the evaluation on legal grounds, our proposed framework achieved performance improvements of 8.6% and 9.1% in the two experimental settings, respectively. In contrast, GPT-4’s performance in the first and second instance settings only reach 13.6% and 22.8%, respectively. This not only indicates significant shortcomings in the capabilities of LLMs in sourcing legal provisions, but also reflects the high challenge of our benchmark. In terms of judgment results evaluation, while all models performed well in the conviction of criminal cases, there is still a significant gap in determining prison term and fines compared to standard results. Furthermore, although the analysis of these systems has shown a certain degree of logicality, there is still room for improvement in terms of correctness and concision.
@@ 5.5 Discussion and Analysis
Judicial Knowledge of LLMs  As indicated in Figure 5 and presented in Table 7 in Appendix, the LLMs demonstrates a high accuracy rate of 96.6% in predicting case types based on limited basic case information. However, its performance in predicting causes of action is less impressive, with only a 35.0% accuracy rate, which increased to just 66.6% even when provided with a list of potential causes. The F1 score of legal grounds generated by LLMs is lower, only 13.6%. This highlights the limitations of LLMs in judicial knowledge.
Figure 5: Judicial knowledge evaluation of LLMs.
Difficulty of Distinct Types of Cases  Table 4 presents the results of our framework in generating legal grounds across different types of cases in the first instance setting. The agent system produces more reliable legal grounds in criminal cases, while its use and understanding of relevant legal statutes in civil and administrative cases are notably weaker. This observation may be attributed to the fact that in criminal cases, the nature of the offense is typically clearer, allowing the agent to more easily apply relevant legal statutes to specific situations. In contrast, civil and administrative cases often involve more complex issues, with multiple vested interests, such as contract disputes, family matters, or government decisions, requiring a deeper understanding of legal and social knowledge.
Figure 6: The absolute difference change.
@@
Table 4: Legal grounds evaluation of AgentCourt.
Judicial knowledge base  With the support of an external knowledge base, the performance of agent system in judicial reasoning improved significantly, with an increase of up to 6.2%. The achievements are also attributed to our designed automatic retrieval module. As shown in Table 6 in Appendix E, through the rough retrieval, the most similar cases only have a 62% consistency in the cause of action with the current cases. However, after the documents re-ranking, the consistency of the cause of action between retrieved cases and the current cases increased to 85%. This improvement proves the effectiveness of our retrieval module.
@@ 6 Conclusion
We introduce SimuCourt, a judicial benchmark to evaluate the judicial analysis and decision-making power of agents. Furthermore, we propose a novel multi-agent framework AgentsCourt, which can sequentially simulate court debate, retrieve precedents, analyze cases, provide legal grounds, and deliver clear judgment. Then, we perform experiments to analyze different modules. The new judicial paradigm we presented effectively simulates the judicial decision-making with multi-agent, which significantly enhances judicial efficiency.
@@ 7 Limitation
  • Our data only includes Chinese documents from "China Judgments Online." Despite our framewok AgentCourt not being specifically designed for the civil law system, testing the agent system with real data from different legal systems is important.
@@
We look forward to further exploring the potential of the judicial knowledge base in future studies.
Appendix A List of Information
The Detailed List is presented in Table 8
Appendix B Data Quality Inspection
@@
Manual Inspection: After completing data annotation and handling private information, we manually inspect the quality of SimuCourt: (1) Case Meeting Standards. The selected samples need to include clear case analysis and facts and have not been overturned in the appellate stage. (2) Accurate Information Annotation. Annotation should ensure the accurate and error-free extraction of key information from the original legal documents, including case analysis, legal grounds, and judgement. (3) Privacy Information Security.In order to safeguard individual privacy and security, it is crucial to ensure that each data entry does not contain any content that could potentially disclose sensitive information about the parties involved. We employ three graduate students to manually review all 420 annotated cases. By carefully scrutinizing, our dataset exhibits a high level of quality. Specific quality metrics and analysis results are shown in Table 5.
Appendix C Data analysis
The cause of action of civil cases statistics in 2022 is shown in Figure 7
@@ Appendix C Data analysis
Precedents Rough retrieval + Re-ranking
Top1 62% 85%
Top2 60% 82%
Top3 61% 80%
Table 6: Cause of action matching
Model Type Pred. Cause Classif. Cause Pred. Legal grounds
GPT-3.5 96.0 20.8 42.5 11.7
GPT-4 97.6 35.0 66.6 13.6
Table 7: Judicial knowledge evaluation of LLMs
Appendix D Data example
We list several cases of criminal, civil and administrative in Table 9-11.
Appendix E Retrieval module
@@
Case Category Case Category
@@
Complaint/Indictment Petition for appeal
Statement of the plaintiff Statement of the appellant
Statement of the defendant/the accused Statement of the appellee
@@
Judicial analysis in the first Instance
legal grounds of the first Instance
Judgement of the first Instance
@@
The court holds that the accused, John Doe, has repeatedly stolen citizens’ property, constituting theft, and should be severely punished. The charges brought by the prosecutor’s office are established. After being apprehended, the accused truthfully confessed to his crimes, voluntarily pleaded guilty, and returned part of the stolen goods, thus is eligible for a lighter punishment according to law. The defense attorney’s reasonable plea for leniency for the accused is accepted.
Legal grounds
@@
Article 64 of the Criminal Law of the People’s Republic of China; Paragraph 3 of Article 67 of the Criminal Law of the People’s Republic of China; Article 264 of the Criminal Law of the People’s Republic of China; Article 15 of the Criminal Procedure Law of the People’s Republic of China.
Judgement
@@
Charge: The defendant is convicted of theft; Prison term: Sentenced to three years and eight months in prison; Fine: Fined ten thousand yuan.
Table 9: Criminal case example.
Cause of action Item Content
Private lending dispute
Case analysis
The court holds that legal private lending is protected by law. The mortgage loan contract between the plaintiff and defendant is lawful and valid, obliging all parties to fully comply. After the plaintiff lent the money, the defendant mortgaged their property as collateral and registered this mortgage. The defendant must repay the principal and interest as agreed or bear the breach of contract responsibilities, including the plaintiff’s legal fees and preservation guarantee fees incurred for debt collection. The court supports the plaintiff’s claim for legal and preservation fees, as stipulated in the contract and evidenced by corresponding receipts.
Legal grounds
Article 389 of the Civil Code of the People’s Republic of China; Article 394 of the Civil Code of the People’s Republic of China; Article 395 of the Civil Code of the People’s Republic of China; Article 400 of the Civil Code of the People’s Republic of China; Article 407 of the Civil Code of the People’s Republic of China; Paragraph 1 of Article 509 of the Civil Code of the People’s Republic of China; Article 577 of the Civil Code of the People’s Republic of China; Article 675 of the Civil Code of the People’s Republic of China; Article 676 of the Civil Code of the People’s Republic of China; Article 67 of the Civil Procedure Law of the People’s Republic of China.
Judgement

### v3

@@ 4.2 Legal Knowledge Base Construction
Journal Articles Journal articles, typically authored by legal experts, can provide in-depth analysis and unique perspectives on specific legal issues. We collect highly-cited journal articles from 2010 to 2023 from the Chinese Legal Resources Knowledge Database  [^4] https://lawnew.cnki.net/.
Precedents We collect all judgement documents of criminal, civil and administrative cases from the China Judgements Online for the years 2017 to 2022. However, as illustrated in Figure 9 in the Appendix, the data exhibits a significant long-tail distribution. To balance the type of case, we limit the number of cases for each cause of action to no more than 20k. For those causes of action with more cases, we retain only the top 20k cases with the longest text as representatives of complex cases.
@@
Table 4: Statistics of our legal knowledge base.
@@ 5 Experiments
Model Legal Articles Judgement Results Case Analysis
@@
5.1 Automatic Evaluation
As example data illustrated in Table 11, the legal articles and judgement are concise and structured. Therefore, we propose corresponding metrics for legal articles and judgement evaluation.
Legal Articles Evaluation The correct legal articles is crucial for a fair judgment. Thus, we employ the strict matching method to assess the legal articles generated by the agent system. Specifically, we calculate the number of entries that match and do not match between the legal articles list of the agent system and the reference legal articles list. These counts are then micro-averaged to determine the overall precision, recall and F1 scores. Details can be found in Table 13.
Judgement Evaluation for Civil and Administrative Cases The judgment of each civil or administrative case may encompass multiple results. While each result typically revolves around a single key point, it may involve specific monetary amounts and interest rate information. Consequently, traditional text matching methods based on similarity struggle to accurately capture these key points. Thus, we employ GPT-4 as an evaluator. Specifically, we separately count the number of matching and non-matching key points in the agent system’s judgment results compared to the reference judgment results. The micro-averaged counts are used to calculate the overall precision, recall and F1 scores. Details is presented in Table 14.
@@ 5.2 Human Evaluation
The case analysis entails intricate logical reasoning and ethical considerations that are challenging to evaluate through automatic metrics or GPT-4. For each setting, we present a panel of three graduate students majoring in law a random sample of 100 entries from each setting and the following binary True/False criteria guidelines: 1) Correctness: Mark true if and only if the analysis is satisfying and considers all parties involved. 2) Logicality: Mark false if the analysis contains any illogical or untrue reasoning. 3) Concision: Mark true if the analysis covers all necessary information without any extra information.
Model Legal Articles Judgement Results
Civil and Admini. Charge Prison term Fine
SimuCourt 0.203 0.507 0.887 0.337 0.500
   w/o Court simulation 0.171 0.473 0.875 0.300 0.462
   w/o Knowledge base 0.145 0.462 0.850 0.312 0.475
   w/o Web search 0.196 0.488 0.865 0.325 0.487
Table 6: Ablation study of our framework in the first instance setting.
@@ 5.3 Baselines
LaWGPT Song et al. (2023)  This is currently the most popular Chinese legal large language model [^5] https://github.com/pengxiao-song/LaWGPT, which has undergone extensive pre-training on Chinese legal corpora and fine-tuning on legal instructions, based on the general Chinese foundation model (Chinese-LLaMA-7B). It possesses strong capabilities in understanding and generating legal content.
@@ 5.4 Main Results
As shown in Table 5, our framework outperforms other models in all aspects. For the evaluation on legal articles, our proposed framework achieved performance improvements of 8.6% and 9.1% in the two experimental settings, respectively. In contrast, GPT-4’s performance in the first and second instance settings only reach 13.6% and 22.8%, respectively. This not only indicates significant shortcomings in the capabilities of LLMs in sourcing legal provisions, but also reflects the high challenge of our benchmark. In terms of judgment results evaluation, while all models performed well in the conviction of criminal cases, there is still a significant gap in determining prison term and fines compared to standard results. Furthermore, although the analysis of these systems has shown a certain degree of logicality, there is still room for improvement in terms of correctness and concision.
@@ 5.5 Discussion and Analysis
Legal Knowledge of LLMs  As indicated in Figure 5, all three language models exhibit excellent performance on the simple task of predicting case types. However, their performance is less impressive on the challenging task of predicting case reasons, the GPT-4 model achieves only 35.4% accuracy, while LaWGPT, which has undergone extensive pre-training with professional knowledge, achieves only 43.7%. For the task of article generation, the performance of all models is poor, with LaWGPT sometimes producing garbled output, resulting in even worse performance.
Figure 5: Legal knowledge evaluation of LLMs.
Difficulty of Distinct Types of Cases  Table 7 presents the results of our framework in generating legal articles across different types of cases in the first instance setting. The agent system produces more reliable legal articles in criminal cases, while its use and understanding of relevant legal statutes in civil and administrative cases are notably weaker. This observation may be attributed to the fact that the civil and administrative cases involve more complex issues, with multiple vested interests, such as contract disputes, family matters, or government decisions, requiring a deeper understanding of legal and social knowledge.
@@
Table 7: Legal articles evaluation of AgentsCourt.
Legal knowledge base  With the support of an external knowledge base, the performance of agent system in judicial reasoning improved significantly, with an increase of up to 6.2%. The achievements are also attributed to our designed automatic retrieval module. As shown in Table 8 in Appendix A, through the rough retrieval, the most similar cases only have a 62% consistency in the cause of action with the current cases. However, after the documents re-ranking, the consistency of the cause of action between retrieved cases and the current cases increased to 85%. This improvement proves the effectiveness of our retrieval module.
@@ 6 Conclusion
We propose a novel multi-agent framework AgentsCourt, which can sequentially simulate court debate, retrieve precedents, analyze cases, provide legal articles, and deliver clear judgment. Furthermore, we introduce SimuCourt, a judicial benchmark to evaluate the judicial analysis and decision-making power of agents. Then, we perform experiments to analyze different modules. The new judicial paradigm we presented effectively simulates the judicial decision-making with multi-agent, which significantly enhances judicial efficiency.
@@ 7 Limitation
  • Our data only includes Chinese documents from "China Judgments Online." Despite our framewok AgentsCourt not being specifically designed for the civil law system, testing the agent system with real data from different legal systems is important.
@@
We look forward to further exploring the potential of the legal knowledge base in future studies.
Appendix A Retrieval Module
Appendix B Example of Court Transcript
We present an example of court transcript simulated by multi-agent debate in Table 12.
Appendix C Data Demonstration
The detailed list is presented in Table 10. Furthermore, we show examples of the first-instance stage in Figure 7 and second-instance stage in Figure 8, respectively.
Appendix D Data Analysis
D.1 Data Description
Our choice of cases is driven by three reasons: (1) Diversity of causes of action. Based on our statistical analysis of data from the China Judgements Online over the past few years, we observed a significant long-tail distribution in various types of cases. For example, as shown in Figure 9, in the total civil cases of 2022, the top 15 causes of action accounted for 66% of the total number of cases. To reflect a broader spectrum of legal practice, we focus on maintaining diversity in the types of causes of action; (2) Clarity of case analysis and facts. We have meticulously selected judgement documents that provide detailed case analysis and clear determine facts for annotation. This aim is to enhance the quality and accuracy of data annotation while aiding agents in better understanding the judicial reasoning and legal articles; (3) Uniqueness and accuracy of judgements. We prioritize cases that are not overturned in appellate review. This ensures the consistency of our evaluation, as these cases have already undergone a rigorous litigation process and the judgements are fair.
D.2 Data Quality Inspection
@@
Manual Inspection: After completing data annotation and handling private information, we manually inspect the quality of SimuCourt: (1) Case Meeting Standards. The selected samples need to include clear case analysis and facts and have not been overturned in the appellate stage. (2) Accurate Information Annotation. Annotation should ensure the accurate and error-free extraction of key information from the original legal documents, including case analysis, legal articles, and judgement. (3) Privacy Information Security.In order to safeguard individual privacy and security, it is crucial to ensure that each data entry does not contain any content that could potentially disclose sensitive information about the parties involved. We employ three graduate students to manually review all 420 annotated cases. By carefully scrutinizing, our dataset exhibits a high level of quality. Specific quality metrics and analysis results are shown in Table 9.
Precedents Rough retrieval + Re-ranking
Top1 62% 85%
Top2 60% 82%
Top3 61% 80%
Table 8: Cause of action matching
@@
Figure 6: The absolute difference change.
Figure 7: An example case of first-instance stage (translated from Chinese).
Figure 8: An example case of second-instance stage (translated from Chinese).
@@
Case type Case type
@@
Indictment Petition for appeal
The point of defense lawyer The point of the appellant
The point of the defendant The point of the appellee
@@
Case analysis Judicial analysis in the first instance
Legal Articles Legal articles of the first instance
Judgement Judgement of the first instance
@@
Theft Case analysis The court holds that the accused, John Doe, has repeatedly stolen citizens’ property, constituting theft, and should be severely punished. The charges brought by the prosecutor’s office are established. After being apprehended, the accused truthfully confessed to his crimes, voluntarily pleaded guilty, and returned part of the stolen goods, thus is eligible for a lighter punishment according to law. The defense attorney’s reasonable plea for leniency for the accused is accepted.
@@
Legal articles Article 64 of the Criminal Law of the People’s Republic of China; Paragraph 3 of Article 67 of the Criminal Law of the People’s Republic of China; Article 264 of the Criminal Law of the People’s Republic of China; Article 15 of the Criminal Procedure Law of the People’s Republic of China.
@@
Judgement Charge: The defendant is convicted of theft; Prison term: Sentenced to three years and eight months in prison; Fine: Fined ten thousand yuan.
Table 11: An example output of criminal case.
Role Record
Judge (1)Determine facts: On October 10, 2022, the defendant John Smith noticed the victim, David Lee, wearing a pair of gold earrings near a hotel. He then followed David Lee into an alley, and taking advantage of David Lee’s unawareness, he snatched the pair of gold earrings from behind. After successfully obtaining them, the defendant John Smith fled the scene… (2) Background of the defendant: John Smith was sentenced to six months’ imprisonment by a certain People’s Court in June 2009 for the crime of theft; John Smith was sentenced to two years’ imprisonment by a certain People’s Court in December 2010 for the crime of robbery… (3) Court Discipline…The plaintiff and the defendant may now make their statements.
Plaintiff The defendant, John Smith, committed the crime of robbery with the purpose of unlawfully obtaining property from others, involving a considerable amount. His actions violate Article 267 of the Criminal Law of the People’s Republic of China. The criminal facts are clear, and the evidence is indeed substantial and sufficient. Therefore, he should be held criminally liable for the crime of robbery.
Defendant I admit that the robbery occurred on October 10, 2022, which was an act of impulse. After the incident, I have truthfully confessed my crime. I hope the judge can give me a lenient punishment.
Plaintiff The defendant has previously been sentenced to a term of imprisonment for intentionally committing a crime. If he commits another crime within five years of completing the sentence, he should be sentenced to a term of imprisonment exceeding the original sentence, which constitutes recidivism.
Defendant I deeply acknowledge that my actions have caused adverse effects on society and irreversible losses to the victims. However, I assure the court that I will actively compensate the victims for their losses. I sincerely repent and am willing to accept the court’s judgment.
Table 12: An example of court transcript simulated by multi-agent debate (translated from Chinese).
Appendix E Details of Automatic Evaluation
E.1 Legal Articles Evaluation
We start by using pattern matching to parse the free text, followed by a hard match against specific legal provisions. For example, as shown in Table 13. Then, with TP (True Positives) = 2, FP (False Positives) = 1, FN (False Negatives) = 2, the corresponding Precision = 2/3, and Recall = 2/4.
E.2 Judgement Evaluation for Civil and Administrative Cases
We utilize GPT-4 to assess the judgment results generated by the model in civil and administrative cases. As shown in Table x, we present an evaluation example, which is also a prompt demonstration for GPT-4.
Reference legal articles Generated legal articles
Article 67, Section 1 of the Criminal Law of the People’s Republic of China Article 67 of the Criminal Law of the People’s Republic of China
Article 52 of Criminal Law of the People’s Republic of China Article 53 of the Criminal Law of the People’s Republic of China
Article 53 of the Criminal Law of the People’s Republic of China Article 52 of Criminal Law of the People’s Republic of China
Article 15 of the Criminal Procedure Law of the People’s Republic of China
Table 13: A comparison between the reference legal articles and the legal articles generated by the model (translated from Chinese).
Reference judgment Generated judgment